### v3

@@ Abstract
Data-hungry HSI classification methods require high-quality labeled HSIs, which are often costly to obtain. This characteristic limits the performance potential of data-driven methods when dealing with limited annotated samples. Bridging the domain gap between data acquired from different sensors allows us to utilize abundant labeled data across sensors to break this bottleneck. In this paper, we propose a novel Attention-Gated Tuning (AGT) strategy and a triplet-structured transformer model, Tri-Former, to address this issue. The AGT strategy serves as a bridge, allowing us to leverage existing labeled HSI datasets, even RGB datasets to enhance the performance on new HSI datasets with limited samples. Instead of inserting additional parameters inside the basic model, we train a lightweight auxiliary branch that takes intermediate features as input from the basic model and makes predictions. The proposed AGT resolves conflicts between heterogeneous and even cross-modal data by suppressing the disturbing information and enhances the useful information through a soft gate. Additionally, we introduce Tri-Former, a triplet-structured transformer with a spectral-spatial separation design that enhances parameter utilization and computational efficiency, enabling easier and flexible fine-tuning. Comparison experiments conducted on three representative HSI datasets captured by different sensors demonstrate the proposed Tri-Former achieves better performance compared to several state-of-the-art methods. Homologous, heterologous and cross-modal tuning experiments verified the effectiveness of the proposed AGT. Code has been released at: https://github.com/Cecilia-xue/AGT.
keywords:
[…]
1 Introduction
@@
Hyperpectral image (HSI) contains rich spectral information about the composition and properties of objects, resulting in a wide range of applications in earth observation, e.g., mineral exploration Carrino et al. (2018), plant stress detection Behmann et al. (2014), and environmental science Transon et al. (2018), etc. Within these applications, HSI classification technology identifies each pixel in an image and assign it to a specific land-cover category. Over the past decade, deep learning (DL) based methods Lin et al. (2013); Chen et al. (2014, 2015); Zhang et al. (2017); Li et al. (2017); Chen et al. (2016); Zhang et al. (2021) dominate the field of HSI classification. From 2013 until now, various optimized convolutional neural networks (ConvNets) have been proposed for HSI classification Zhang et al. (2017); Li et al. (2017); Chen et al. (2016). Among them, 3D ConvNet-based HSI classification methods have become the mainstream in this field, because the 3D convolution is inherently well-suited for processing the HSI with 3D structure. Recently, the emergence of the vision transformer (ViT) has disrupted this situation and researchers have begun exploring how to apply ViT models to handle the HSI classification task. Hong et al. Hong et al. (2021) were the first to apply pure transformers to HSI classification. In the same trend, impressive results have been achieved by combining the transformer with ConvNet Sun et al. (2022), architecture search (NAS) algorithm Xue et al. (2022) and new absolute position embedding Jiang et al. (2024).
@@
The performance of aforementioned data-hungry methods are closely related to the quantity and quality of labeled data. However, annotating HSIs requires field surveys by professionals, making the process time-consuming and expensive. The number of labeled hyperspectral data is relatively small. To alleviate this problem, some early work Zhang et al. (2019); Yang et al. (2017) explores how to use additional labelled data from other domains to further push back these limits. Two-CNN-transfer Yang et al. (2017) first investigates transfer learning between the data from the same sensor in HSI classification. On similar trend,  Zhang et al. (2019) concentrates on exploiting data from other sensors and transferring the knowledge. Such previous work Zhang et al. (2019); Yang et al. (2017) both bridge the domain gap between source and target domains by fine tuning the ConvNets. However, such fine tuning methods do not perform effectively enough when meeting the transformer architecture. Because ConvNets optimize convolutional filter parameters during training, while transformer architectures learn specific pixel relationships within the training data. Compared with ConvNets, the transformer-based models are always more sensitive to the characteristics of the training data. Given these differing learning approaches, the fine tuning transfer learning strategies proposed for ConvNet-based methods are not able to be directly applied to the transformer framework. Although some recent approaches Gao et al. (2021); Hu et al. (2022); Sung et al. (2022) have been proposed to the transformer architectures, they cannot learn domain-invariant features effectively on HSIs. Because HSI datasets acquired by various sensors display distinct structural characteristics. Tuning methods designed for unimodal RGB images cannot takes advantage of the cross-modal discrepancies and handle the gaps between cross-sensor HSIs well.
@@
    A novel Attention-Gated Tuning (AGT) strategy is proposed for transformer-based HSI classification methods to address conflicts in cross-sensor and even cross-modality data.
@@
The remaining sections of this paper are structured as follows. Section 2 presents a review of related work. In Section 3 , we provide a detailed explanation of our proposed Tri-Former model and AGT strategy. Algorithm implementation details are outlined in Section 4 , followed by an extensive evaluation and comparison with state-of-the-art competitors. Finally, in Section 5, we conclude this work.
2 Related Work
2.1 ConvNets for Hyperspectral Image Classification
@@
ConvNets Krizhevsky et al. (2012); Simonyan and Zisserman (2014); Szegedy et al. (2017); He et al. (2016); Huang et al. (2017); Howard et al. (2017); Liu et al. (2022) emerged as a powerful tool for visual classification, revolutionizing the field by shifting focus from handcrafted features to data-driven architecture design. The adeptness of convolutional networks in feature extraction perfectly aligns with the demands of hyperspectral image classification. The emergence of convolutional networks has also revolutionized the field of HSI classification, elevating the accuracy of classification to a new level.
@@
Over the past decade, various convolutional variants have been proposed for HSI classification. 1D ConvNet and 2D ConvNet are first introduced into HSI classification. Mei and Ling Mei et al. (2016); Zhang and Li (2016) employed 1D ConvNets to perform convolution along the spectral dimension, effectively extracting spectral features. Meanwhile, a series of HSI classification approaches Makantasis et al. (2015); Yue et al. (2015) based on 2D ConvNet have been proposed. 1D ConvNet considers spectral dimension, while 2D ConvNet always focuses on spatial dimension. Later, researchers pursued dual-channel ConvNet structures after recognizing the potential of combining spectral and spatial information. Zhang et al. Zhang et al. (2017) and Yang et al.  Yang et al. (2017) introduced dual-channel ConvNet models that effectively integrated 1D ConvNet and 2D ConvNet, achieving further improvements in accuracy for HSI classification. Since 2016, 3D ConvNet based methods have become the mainstream as 3D convolution naturally extracts spectral-spatial features from 3D HSIs. Li et al. Li et al. (2017) and Chen et al. Chen et al. (2016) pioneered this approach, constructing 3D ConvNet architectures to capture spatial-spectral patterns efficiently. The inherent 3D nature of HSIs was leveraged, enabling comprehensive feature extraction and improving classification performance. After that, researchers mainly focus on designing HSI characteristic-aligned 3D ConvNets. Efficient residual structures were introduced by Zhong et al. Zhong et al. (2018), integrating spectral and spatial residual modules to improve feature learning. Zhang et al. Zhang et al. (2019) focused on lightweight 3D-CNN designs and transfer learning strategies to handle small sample problems. Getting inspiration from DARTS, Chen et al. Chen et al. (2019) proposed a 3D Auto-CNN for HSI classification. In the preprocessing stage, 3D Auto-CNN heavily compresses the spectral dimension of raw HSIs through point wise convolution. Under this basis, 3DANAS Zhang et al. (2021) and HyT-NAS Xue et al. (2022) are designed, which introduce the NAS algorithm into an efficient 3D pixel-to-pixel classification framework.
@@
2.2 Vision Transformers for Hyperspectral Image Classification
Transformer model was initially introduced for Natural Language Processing (NLP) by Vaswani et al. Vaswani et al. (2017). It brought about a revolutionary change in processing sequential data through its attention mechanism. In the field of computer vision, ViT Dosovitskiy et al. (2020) builds on the success of the transformer, which was originally designed for NLP.
More recently, within the HSI classification domain, researchers have redirected their focus towards the transformer model in their quest for enhanced performance Hong et al. (2021); He et al. (2021); Qing et al. (2021). For instance, He et al. He et al. (2021) introduced a spatial-spectral transformer that combines a ConvNet for capturing spatial information with a ViT for extracting spectral relationships. Similarly, the method proposed in Qing et al. (2021) adopts a spectral relationship extraction transformer along with several decoders. SpectralFormer Hong et al. (2021) mines and represents the sequence attributes of spectral signatures well from input HSIs using the transformer model. It extracts spectrally local sequence information from neighboring bands of HSIs. Meanwhile, HyT-NAS Xue et al. (2022) first combines the NAS and transformer for handling the HSI classification task. An emerging transformer module is grafted on the automatically designed ConvNet to add global information to local region focused features. The searching space in HyT-NAS is hybrid, suitable for the HSI data with a relatively low spatial resolution and an extremely high spectral resolution. To better utilize high-level semantic spectral-spatial features, a spectral-spatial feature tokenization transformer (SSFTT) method Sun et al. (2022) is proposed. Specifically, a spectral–spatial feature extraction module is built to extract low-level shallow features and a gaussian weighted feature tokenizer is introduced for feature transformation. Based on this, the transformed features are input into the transformer encoder module for stronger feature representation and learning. Another model GraphGST Jiang et al. (2024) incorporates a new absolute positional encoding (APE) to capture pixel positional sequences and uses self-supervised learning for cross-view contrastive learning, improving representation and capturing local-to-global correlations in HSI classification tasks.
@@ 2.2 Vision Transformers for Hyperspectral Image Classification
2.3 Fine Tuning Methods
In real-world scenarios, obtaining a sufficient amount of training data for HSI classification is challenging due to the expensive sensor cost and labeling cost. Therefore, designing models with low training sample requirements has consistently been a goal within the HSI classification domain. A classical approach to addressing this challenge is to pretrain the method on the sufficiently labeled data (maybe from another domain) first, and then transfer the acquired knowledge and representations to the target domain Pan and Yang (2009).
Fine Tuning ConvNets in HSI classification. As a pioneering effort to address HSI classification problems with limited training samples, Yang et al.  Yang et al. (2017) introduced the transfer learning strategy into a two-branch ConvNet to extract more relevant features from HSIs. In Zhang et al. (2019), Zhang et al. further improved this idea and transferred feature extraction capacity among heterogeneous HSI datsets and even cross-modal datasets. Unfortunately, these methods may not exhibit effective performance on transformer architecture. In LWNet Zhang et al. (2019) and Two-CNN Yang et al. (2017), fine tuning strategies are employed on ConvNet architectures, yielding promising results. However, unlike ConvNets, which learn convolutional filter parameters from extensive datasets, transformer-based methods learn specific pixel relationships within the training data. Given these differing learning approaches, the fine tuning transfer learning strategy used in ConvNet-based methods exhibit limited effectiveness in the transformer framework.
Full-Parameter Fine Tuning. Due to the rapid advancements in vision language models, an increasing number of fine tuning methods targeting transformer architectures have been proposed. As shown in Figure 1(a), Full Parameter Fine Tuning (FPFT) adjusts all layers and parameters based on the pre-trained model to adapt it to a specific task. This process usually uses a small learning rate and task-specific data to fully utilize the general features of the pre-trained model, but may require more computing resources.
@@ 2.3 Fine Tuning Methods
Parameter-Efficient Fine Tuning. Parameter-Efficient Fine Tuning (PEFT) technology aims to improve the performance of pre-trained models on new tasks by minimizing the number of fine tuning parameters and computational complexity. In this way, even if computing resources are limited, the knowledge of the pre-trained model can be used to quickly adapt to new tasks and achieve efficient transfer learning. Therefore, PEFT technology can greatly shorten the model training time and computational cost while improving the model effect, allowing more people to participate in deep learning research. A popular line of work on PEFT is to add a few trainable parameters without changing the original parameters of the pre-trained model. Figure 1(b)-Figure 1(e) illustrate the information forward and gradient backpropagation process of some representative PEFT approaches. Adapter tuning Gao et al. (2021) approaches always insert small neural network modules, called "adapters", between each layer of the model or between certain specific layers. These adapters are trainable, while the parameters of the original model remain unchanged. Similar to Adapter tuning, LoRA Hu et al. (2022) adds small, low-rank matrices to the key layers of the model to adjust the model’s behavior. Prompt tuning Lester et al. (2021) adds learnable embedding vectors as prompts to the input of the pre-trained language model. These prompts are designed to be updated during training to guide the model to output more useful responses for specific tasks. Without inserting additional parameters inside backbone networks, another line of work on PEFT is to add an additional small and separate network. Zhang et al. Zhang et al. (2020) proposed Side-tuning, which adapts a pre-trained network by training a lightweight “side” network that is fused with the (unchanged) pre-trained network via summation. In a similar trend, LST Sung et al. (2022) trains a ladder side network. This small and separate network takes intermediate activations as input via shortcut connections (called ladders) from backbone networks and makes predictions.
@@
In this paper, in order to make full use of labeled cross-sensor data, we propose an attention-gated tuning (AGT) strategy and employ it to fintune the proposed efficient Tri-Former. The proposed approach demonstrates effectiveness across homogeneous, heterogeneous, and even cross-modality cases. Although the proposed AGT takes inspiration and has a similar parallel structure to LST, we argue that there are major differences in motivations, key problems and architecture designs between the two methods. Detail comparisons and comprehensive analysis are presented in Section 4.3. The corresponding experimental results are outlined in Section 4.6-4.8.
3 Proposed Method
In this section, we provide a detailed introduction to the proposed Tri-Former model and AGT strategy. We start by giving a concise overview of the vanilla visual transformer architecture before diving into the architecture of our Tri-Former model. Following that, we elaborate on the details of the AGT strategy, emphasizing its differences from previous approaches.
3.1 Primary introduction of ViT model

### v4

@@ Abstract
Data-driven HSI classification methods require high-quality labeled HSIs, which are often costly to obtain. This characteristic limits the performance potential of data-driven methods when dealing with limited annotated samples. Bridging the domain gap between data acquired from different sensors allows us to utilize abundant labeled data across sensors to break this bottleneck. In this paper, we propose a novel Attention-Gated Tuning (AGT) strategy and a triplet-structured transformer model, Tri-Former, to address this issue. The AGT strategy serves as a bridge, allowing us to leverage existing labeled HSI datasets, even RGB datasets to enhance the performance on new HSI datasets with limited samples. Instead of inserting additional parameters inside the basic model, we train a lightweight auxiliary branch that takes intermediate features as input from the basic model and makes predictions. The proposed AGT resolves conflicts between heterogeneous and even cross-modal data by suppressing the disturbing information and enhances the useful information through a soft gate. Additionally, we introduce Tri-Former, a triplet-structured transformer with a spectral-spatial separation design that enhances parameter utilization and computational efficiency, enabling easier and flexible fine-tuning. Comparison experiments conducted on three representative HSI datasets captured by different sensors demonstrate the proposed Tri-Former achieves better performance compared to several state-of-the-art methods. Homologous, heterologous and cross-modal tuning experiments verified the effectiveness of the proposed AGT. Code has been released at: https://github.com/Cecilia-xue/AGT.
Index Terms:
hyperspectral image classification, cross-sensor tuning, attention gated tuning, vision transformer, triplet-structured transformer
I Introduction
@@
Hyperpectral image (HSI) contains rich spectral information about the composition and properties of objects, resulting in a wide range of applications in earth observation, e.g., mineral exploration [1], plant stress detection [2], and environmental science [3], etc. Within these applications, HSI classification technology identifies each pixel in an image and assign it to a specific land-cover category. Over the past decade, deep learning (DL) based methods [4, 5, 6, 7, 8, 9, 10] dominate the field of HSI classification. From 2013 until now, various optimized convolutional neural networks (ConvNets) have been proposed for HSI classification [7, 8, 9]. Among them, 3D ConvNet-based HSI classification methods have become the mainstream in this field, because the 3D convolution is inherently well-suited for processing the HSI with 3D structure. Recently, the emergence of the vision transformer (ViT) has disrupted this situation and researchers have begun exploring how to apply ViT models to handle the HSI classification task. Hong et al. [11] were the first to apply pure transformers to HSI classification. In the same trend, impressive results have been achieved by combining the transformer with ConvNet [12], architecture search (NAS) algorithm [13] and new absolute position embedding [14].
@@
The performance of aforementioned data-hungry methods are closely related to the quantity and quality of labeled data. However, annotating HSIs requires field surveys by professionals, making the process time-consuming and expensive. The number of labeled hyperspectral data is relatively small. To alleviate this problem, some early work [15, 16] explores how to use additional labelled data from other domains to further push back these limits. Two-CNN-transfer [16] first investigates transfer learning between the data from the same sensor in HSI classification. On similar trend, method [15] concentrates on exploiting data from other sensors and transferring the knowledge. Such previous work [15, 16] both bridge the domain gap between source and target domains by fine tuning the ConvNets. However, such fine tuning methods do not perform effectively enough when meeting the transformer architecture. Because ConvNets optimize convolutional filter parameters during training, while transformer architectures learn specific pixel relationships within the training data. Compared with ConvNets, the transformer-based models are always more sensitive to the characteristics of the training data. Given these differing learning approaches, the fine tuning transfer learning strategies proposed for ConvNet-based methods are not able to be directly applied to the transformer framework. Although some recent approaches [17, 18, 19] have been proposed to the transformer architectures, they cannot learn domain-invariant features effectively on HSIs. Because HSI datasets acquired by various sensors display distinct structural characteristics. Tuning methods designed for unimodal RGB images cannot takes advantage of the cross-modal discrepancies and handle the gaps between cross-sensor HSIs well.
@@
    A novel Attention-Gated Tuning (AGT) strategy is proposed for transformer-based HSI classification methods. To the best of our knowledge, in HSI classification field, this is the first tuning strategy that leverages transformer structures for cross-sensor and cross-modal adaptation.
@@
In summary, while HSIs provide rich and detailed spectral information, their complex structure necessitates advanced deep learning models for robust feature extraction. Transformer-based architectures have demonstrated outstanding performance in this regard. However, their high data requirements pose significant challenges given the limited availability of annotated HSIs. Motivated by this gap, our approach leverages transfer learning by integrating the powerful feature extraction capabilities of transformers with our novel AGT strategy and Tri-Former architecture. This integration effectively addresses the small sample problem in HSI classification. The remaining sections of this paper are structured as follows. Section II presents a review of related work. In Section III, we provide a detailed explanation of our proposed Tri-Former model and AGT strategy. Algorithm implementation details are outlined in Section IV, followed by an extensive evaluation and comparison with state-of-the-art competitors. Finally, in Section V, we conclude this work.
II Related Work
II-A ConvNets for Hyperspectral Image Classification
@@
ConvNets [20, 21, 22, 23, 24, 25, 26] emerged as a powerful tool for visual classification, revolutionizing the field by shifting focus from handcrafted features to data-driven architecture design. The adeptness of convolutional networks in feature extraction perfectly aligns with the demands of hyperspectral image classification. The emergence of convolutional networks has also revolutionized the field of HSI classification, elevating the accuracy of classification to a new level.
@@
Over the past decade, various convolutional variants have been proposed for HSI classification. 1D ConvNet and 2D ConvNet are first introduced into HSI classification. Mei and Ling [27, 28] employed 1D ConvNets to perform convolution along the spectral dimension, effectively extracting spectral features. Meanwhile, a series of HSI classification approaches [29, 30] based on 2D ConvNet have been proposed. 1D ConvNet considers spectral dimension, while 2D ConvNet always focuses on spatial dimension. Later, researchers pursued dual-channel ConvNet structures after recognizing the potential of combining spectral and spatial information. Zhang et al. [7] and Yang et al.  [16] introduced dual-channel ConvNet models that effectively integrated 1D ConvNet and 2D ConvNet, achieving further improvements in accuracy for HSI classification. Since 2016, 3D ConvNet based methods have become the mainstream as 3D convolution naturally extracts spectral-spatial features from 3D HSIs. Li et al. [8] and Chen et al. [9] pioneered this approach, constructing 3D ConvNet architectures to capture spatial-spectral patterns efficiently. The inherent 3D nature of HSIs was leveraged, enabling comprehensive feature extraction and improving classification performance. After that, researchers mainly focus on designing HSI characteristic-aligned 3D ConvNets. Efficient residual structures were introduced by Zhong et al. [31], integrating spectral and spatial residual modules to improve feature learning. Zhang et al. [15] focused on lightweight 3D-CNN designs and transfer learning strategies to handle small sample problems. Getting inspiration from DARTS, Chen et al. [32] proposed a 3D Auto-CNN for HSI classification. In the preprocessing stage, 3D Auto-CNN heavily compresses the spectral dimension of raw HSIs through point wise convolution. Under this basis, 3DANAS [10] and HyT-NAS [13] are designed, which introduce the NAS algorithm into an efficient 3D pixel-to-pixel classification framework.
@@
II-B Vision Transformers for Hyperspectral Image Classification
Transformer model was initially introduced for Natural Language Processing (NLP) by Vaswani et al. [33]. It brought about a revolutionary change in processing sequential data through its attention mechanism. In the field of computer vision, ViT [34] builds on the success of the transformer, which was originally designed for NLP.
More recently, within the HSI classification domain, researchers have redirected their focus towards the transformer model in their quest for enhanced performance [11, 35, 36]. For instance, He et al. [35] introduced a spatial-spectral transformer that combines a ConvNet for capturing spatial information with a ViT for extracting spectral relationships. Similarly, the method proposed in [36] adopts a spectral relationship extraction transformer along with several decoders. SpectralFormer [11] mines and represents the sequence attributes of spectral signatures well from input HSIs using the transformer model. It extracts spectrally local sequence information from neighboring bands of HSIs. Meanwhile, HyT-NAS [13] first combines the NAS and transformer for handling the HSI classification task. An emerging transformer module is grafted on the automatically designed ConvNet to add global information to local region focused features. The searching space in HyT-NAS is hybrid, suitable for the HSI data with a relatively low spatial resolution and an extremely high spectral resolution. To better utilize high-level semantic spectral-spatial features, a spectral-spatial feature tokenization transformer (SSFTT) method [12] is proposed. Specifically, a spectral–spatial feature extraction module is built to extract low-level shallow features and a gaussian weighted feature tokenizer is introduced for feature transformation. Based on this, the transformed features are input into the transformer encoder module for stronger feature representation and learning. Another model GraphGST [14] incorporates a new absolute positional encoding (APE) to capture pixel positional sequences and uses self-supervised learning for cross-view contrastive learning, improving representation and capturing local-to-global correlations in HSI classification tasks.
@@ II-B Vision Transformers for Hyperspectral Image Classification
II-C Fine Tuning Methods
In real-world scenarios, obtaining a sufficient amount of training data for HSI classification is challenging due to the expensive sensor cost and labeling cost. Therefore, designing models with low training sample requirements has consistently been a goal within the HSI classification domain. A classical approach to addressing this challenge is to pretrain the method on the sufficiently labeled data (maybe from another domain) first, and then transfer the acquired knowledge and representations to the target domain [37].
Fine Tuning ConvNets in HSI classification. As a pioneering effort to address HSI classification problems with limited training samples, Yang et al.  [16] introduced the transfer learning strategy into a two-branch ConvNet to extract more relevant features from HSIs. In [15], Zhang et al. further improved this idea and transferred feature extraction capacity among heterogeneous HSI datsets and even cross-modal datasets. Unfortunately, these methods may not exhibit effective performance on transformer architecture. In LWNet [15] and Two-CNN [16], fine tuning strategies are employed on ConvNet architectures, yielding promising results. However, unlike ConvNets, which learn convolutional filter parameters from extensive datasets, transformer-based methods learn specific pixel relationships within the training data. Given these differing learning approaches, the fine tuning transfer learning strategy used in ConvNet-based methods exhibit limited effectiveness in the transformer framework.
Full-Parameter Fine Tuning. Due to the rapid advancements in vision language models, an increasing number of fine tuning methods targeting transformer architectures have been proposed. As shown in Fig. 1(a), Full Parameter Fine Tuning (FPFT) adjusts all layers and parameters based on the pre-trained model to adapt it to a specific task. This process usually uses a small learning rate and task-specific data to fully utilize the general features of the pre-trained model, but may require more computing resources.
@@ II-C Fine Tuning Methods
Parameter-Efficient Fine Tuning. Parameter-Efficient Fine Tuning (PEFT) technology aims to improve the performance of pre-trained models on new tasks by minimizing the number of fine tuning parameters and computational complexity. In this way, even if computing resources are limited, the knowledge of the pre-trained model can be used to quickly adapt to new tasks and achieve efficient transfer learning. Therefore, PEFT technology can greatly shorten the model training time and computational cost while improving the model effect, allowing more people to participate in deep learning research. A popular line of work on PEFT is to add a few trainable parameters without changing the original parameters of the pre-trained model. Fig. 1(b)-Fig. 1(e) illustrate the information forward and gradient backpropagation process of some representative PEFT approaches. Adapter tuning [17] approaches always insert small neural network modules, called ”adapters”, between each layer of the model or between certain specific layers. These adapters are trainable, while the parameters of the original model remain unchanged. Similar to Adapter tuning, LoRA [18] adds small, low-rank matrices to the key layers of the model to adjust the model’s behavior. Prompt tuning [38] adds learnable embedding vectors as prompts to the input of the pre-trained language model. These prompts are designed to be updated during training to guide the model to output more useful responses for specific tasks. Without inserting additional parameters inside backbone networks, another line of work on PEFT is to add an additional small and separate network. Zhang et al. [39] proposed Side-tuning, which adapts a pre-trained network by training a lightweight “side” network that is fused with the (unchanged) pre-trained network via summation. In a similar trend, LST [19] trains a ladder side network. This small and separate network takes intermediate activations as input via shortcut connections (called ladders) from backbone networks and makes predictions.
@@
In this paper, in order to make full use of labeled cross-sensor data, we propose an attention-gated tuning (AGT) strategy and employ it to fintune the proposed efficient Tri-Former. The proposed approach demonstrates effectiveness across homogeneous, heterogeneous, and even cross-modality cases. Although the proposed AGT takes inspiration and has a similar parallel structure to LST, we argue that there are major differences in motivations, key problems and architecture designs between the two methods. Detail comparisons and comprehensive analysis are presented in Section IV.C. The corresponding experimental results are outlined in Section IV.J-IV.I.
III Proposed Method
In this section, we provide a detailed introduction to our Tri-Former model and the proposed Attention-Gated Tuning (AGT) strategy. We first briefly revisit the vanilla Vision Transformer (ViT) architecture, then describe the Tri-Former architecture, introducing how the Spectral-Spatial Parallel (SSP) block and 3D convolution reduce redundancy and enhance learning for HSI data. Following that, we elaborate on the details of the AGT strategy, emphasizing its ability to leverage cross-sensor or cross-modality data and differences from previous approaches. Finally, following [40, 41, 42, 43], we provide the pseudo code of the working flow of proposed method in Algorithm 1.
III-A Primary introduction of ViT model